40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 253 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  4370   Wed Mar 2 22:04:22 2011 KevinUpdateElectronicsPOY Shot Noise Measurement

The previous measurement for the shot noise of POY had the dark noise at ~100 nV/rtHz. I redid the measurement and got 26 nV/rtHz for the dark noise. I think that when I made the previous measurement, the spectrum analyzer had automatically added some attenuation to the input that I failed to remove. This added attenuation raised the noise floor of the measurement making the dark noise of POY appear larger than it is.

The updated measurement can be found on the wiki at http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/POY.

  4369   Wed Mar 2 18:08:43 2011 AidanUpdateGreen LockingGhost beams on green

Kiwamu and I noticed that there is a ghost beam on the green beam going into the ETM. What we see is some interference fringes on the edge of the transmission of the green beam through the dichroic beam splitter (DCBS). If we look at the reflection from the dichroic beam splitter these are much more pronounced.

The spacing of the fringes (about 2 per 10mm) indicates an angle between the fields of around 0.1 mrad.

We were able to cause significant motion of the fringes by pushing on the knobs of the steering mirrors that steer the beam into the DCBS. A rough calculation of the derivative of optical path difference between the ghost and the primary beam as a function of input angle gives about 15 microns per mrad. What filtering the effect the arm cavity will have on the ghost beam is not immediately clear, but the numbers shouldn't be too difficult to determine.

 

Attachment 1: Ghost_Beam_at_ETM_DCBS.pdf
Ghost_Beam_at_ETM_DCBS.pdf
  4368   Wed Mar 2 17:19:58 2011 AidanConfigurationGreen LockingMoved PDH PD on end table

As previously noted, there are multiple beams coming back from the ETM surface onto the PDH PD on the end table. They are spread out in a vertical pattern. All the spots swing together (as the ETM moves?).

I moved the PDH Green PD on the end table so that it was further away from the Faraday and I added an iris in between the Faraday and the PD. Now only the principle reflection from the ETM is incident on the PD. See attached photos. In order to sneak past the neighbouring optics, I had to steer the beam down a bit, so the PD is now lower than it previously was.

Just FYI: the angle between the returning beams is about 5 or 6 mrad at the PD. Before the beams get to the PD they go through a telescope that de-magnifies the beam by about 5 or 6 times. This implies that the angle between adjacent returning beams from the ETM is around 1 mrad at the ETM.

This does make the position of the spot on the PD more susceptible to the alignment of the ETM - we should use a short focal length lens and image the ETM plane onto the PD.

 

First image - overview of table

Second image - the three returning beams immediately before the IRIS

Third image - a close up of the IRIS and PDH PD. 

 

 

 

Attachment 1: P1000223-a.jpg
P1000223-a.jpg
Attachment 2: P1000218.jpg
P1000218.jpg
Attachment 3: P1000224.jpg
P1000224.jpg
  4367   Wed Mar 2 16:51:53 2011 steveConfigurationGreen Lockingmech shutter in place at the south end

I moved old POX shutter from ITMY optical table to the south end. MEDM POX mechanical shutter screen is now closing the green beam  injection into the Y arm.

I kluged in a 40m long bnc cable that Alberto left on the floor for control. It is labelled POX-sht  This is a temporary set up.

  4366   Wed Mar 2 04:01:51 2011 KojiUpdateIOOMC servo improvement

[Koji / Rana]

- Since the MC servo had UGF up to ~20kHz and huge servo bump at 50kHz, we needed more phase between 20kHz to 100kHz.

- Today a phase compensation filter in a Pomona box has been inserted between the MC servo box and the FSS box.
  This is a passive filter with zero@14kHz and pole@140kHz. We obtain ~60deg at around 50kHz.

- After the insertion, the lock of the MC was achieved immediately. The overall gain as well as the PZT fast gain was tweaked
  such that the PC feedback is reduced down to 1~2.

- The OLTF has been measured.
  The insertion of the filter change increased the UGF to 130kHz even with "40:4kHz" and double super boost turned on.

  The phase margin is 54deg. Quite healthy.

- Rana modified the existed Auto Locker script.
  It is now continuously running on op340m!
  We made a couple of testsif it correctly relock the MC and it did. VERY COOL.

-----------------

NEXT STEPS
- Measure the PMC cavity pole
- Measure the circuit TF and try to shave off the phase lag.
- Measure the PZT resonance of the NPRO and put notch in the PZT path
- Increase the UGF / measure the openloop TF

Attachment 1: IMG_3904.jpg
IMG_3904.jpg
Attachment 2: MC_OLTF.pdf
MC_OLTF.pdf
  4365   Tue Mar 1 08:42:18 2011 AidanUpdateelogRestarted the elog this morning

 The elog was dead this morning. I reanimated it. It is now undead.

Attachment 1: Zombie.gif
Zombie.gif
  4364   Mon Feb 28 11:22:40 2011 josephbSummaryGeneralto do list
Quote:

- Where is the CDS TO DO ==> Joe

CDS To Do:

1) Get ETMY working - figure out why signals are not getting past the AI board (D000186) to the coils.

2) Get TDS and command line AWG stuff working

3) Get c1ass and new c1lsc (with Koji) fully integrated with the rest of the system.

4) Get CDS software instructions up to date and well organized.

5) Redo cabling  and generally make it a permanent installation instead of hack job:

   a) Measure cable lengths, check connectors, wire with good routes and ensure strain relief.  Ensure proper labeling

   b) Get correct length fiber for c1sus RFM and timing.

   c) Fix up final BO adapter box and DAC boxes.

   d) Make boxes for the AA filter adapters which are currently just hanging.

   e) Get two "faceplates" for the cards in the back of the ETMY IO chassis so they can screwed down properly.

   f) Remove and properly store old, unused cables, boards, and anything else.

6) Create new documentation detailing the current 40m setup, both DCC documents and interactive wiki.

7)  Setup an Ubuntu work station using Keith's wiki instructions

 

Simulated Plant To Do:

1)Create simulated plant to interface with current end mass controls (say scx).

2) Create proper filters for pendulum and noise generation, test suspension.

3) Propagate to all other suspensions.

4) Working on simulated IFO plant to connect to LSC.  Create filters for near locked (assume initial green control perhaps) state.

5) Test LSC controls on simulated IFO.

6) Fix c code so there's seamless switching between simulated and real controls.

CDS Status:

MC damp dataviewer diaggui AWG c1lsc c1ioo c1sus c1iscex c1iscey RFM The Dolphins Sim.Plant Frame builder TDS Cabling
                             
  4363   Sun Feb 27 13:09:56 2011 ranaUpdateGreen Lockingsidebands on beatnote

When Koji and I were massaging the MC, we noticed that the oscillations were at 48.5 kHz. They were pretty huge and are probably what you're seeing on the beat. My guess is that they are the PZT resonances of the PSL 2W NPRO; we need to put a notch in the FSS box - it still has the notch from the old NPRO.

  4362   Sun Feb 27 09:43:59 2011 AidanUpdateGreen Lockingsidebands on beatnote

Can we set up a fiber-PD on the end table to look at the beat between the "end laser IR beam" and the "PSL IR beam fiber-transmitted end beam"? 

We should see the same thing on that PD that we see on the green PD (plus any fiber noise and I'm not really sure how much that'll be off the top of my head). If we unlock the lasers from the arm cavity then the free-running noise of the lasers wrt to each other will probably swamp the 50kHz and 150kHz signals. Maybe we could lock the end laser to the free-running PSL by demodulating the beat note signal from the fiber-PD and then we could look for the extra sidebands in the IN-LOOP signal. Then we could progressively lock the PSL to the MC and arm cavity and see if the sidebands appear on the fiber-PD at some point in that process. 

It's possible that the 216kHz drive of the PZT on the Innolight is somehow driving up some sub-harmonics in the crystal. I think this is unlikely though: if you look at Mott's measurements of the Innolight PZT response, there are no significant PM resonances at 50 or 150kHz.

 

Quote:

Other than the +/-216 kHz sidebands, we can see some funny peaks at +/- 50 kHz and +/-150 kHz.

 

Quote: #4351 by Aidan

So, on the vertex PD, the power of the 80MHz +/-200kHz sidebands should be around sqrt(0.15)*0.05 = 0.02 = 2% of the 80MHz beatnote.

Once we get the green and IR locked to the arm again, we're going to look for the sidebands around the beatnote.

 

  4361   Sat Feb 26 02:33:16 2011 kiwamuUpdateGreen Lockingsidebands on beatnote

The power ratio of the beatnote signal vs. the 216kHz sideband has been measured.

The measured ratio was -55 dB, which is smaller by about 20 dB than Aidan's estimation.

To confirm this fact we should check the modulation depth of the end PDH somehow.

 

The below is a picture showing the sidebands around the beatnote locked at 66.45 MHz.

Other than the +/-216 kHz sidebands, we can see some funny peaks at +/- 50 kHz and +/-150 kHz

I wonder if they come from the servo oscillation of the MC servo whose UGF is at 24 kHz.  We can check it by unlocking the MC.

beat_note.png

Quote: #4351 by Aidan

So, on the vertex PD, the power of the 80MHz +/-200kHz sidebands should be around sqrt(0.15)*0.05 = 0.02 = 2% of the 80MHz beatnote.

Once we get the green and IR locked to the arm again, we're going to look for the sidebands around the beatnote.

  4360   Sat Feb 26 00:25:38 2011 KojiUpdateIOOMC servo improvement

[Rana / Koji]

The MC servo loop has been investigated as the MC servo was not an ideal state.

With the improved situation by us, the attached setting is used for the MC and the FSS.
The current UGF is 24kHz with phase margin is ~15deg, which is unbearably small.
We need to change the phase compensation in the FSS box some time in the next week.


- We found the PD has plenty of 29.5MHz signal in in-lock state. This was fixed by reducing the LO power and the modulation depth.

- The LO power for the MC demodulator was ~6dBm. As this was too high for the demodulator, we have reduced it down to 2dBm
by changing attenuator to 12dB (at 6 oclock of the dial) on the AM stabilization box.

- The RF power on the MC PD was still too high. The PD mush have been saturated. So the modulation slider for 29.5MHz was moved
from 0.0 to 5.0. This reduced the 29.5MHz component. (But eventually Koji restored the modulation depth after the servo shape has been modified.)

- The openloop gain of the loop has been measured using EXC A/TEST1/TEST2. The UGF was ~5kHz with the phase mergin of ~10deg. 

- This quite low phase margin is caused by the fact that the loop has f^-2 shape at around 4k-100kHz. The reference cavity has
the cavity pole of 40kHz or so while the IMC has the pole of ~4kHz. Basically we need phase lead at  around 10-100kHz.

- We decided to turn off (disable) 40:4000 boost of the MC servo to earn some phase. Then MC did not lock. This is because the LF gain was not enough.
So put Kevin's pomona box in the FAST PZT path (1.6:40). By this operation we obtain ~75deg (max) at 560Hz, ~35deg at 5kHz, ~20deg at 10kHz.

- In this setup the UGF is 24kHz. Still the phase margin is ~15kHz. This phase lag might be cause by 1)  the MC servo circut 2) PMC cavity pole

NEXT STEP

- Put/modify phase lead in the FSS box.
- Measure the PMC cavity pole
- Measure and put notch in the PZT path
- Increase the UGF / measure the openloop TF

Attachment 1: fss_servo.png
fss_servo.png
Attachment 2: mc_servo.png
mc_servo.png
  4359   Fri Feb 25 14:50:16 2011 KojiSummaryGeneralto do list

- Put priority on the list

- Put names on the items

- Where is the CDS TO DO ==> Joe

-

- Remote disconnection of the greeen PDH 

- What is the situation of the PD DC for the LSC PDs?

- SUS Satelite box Resister replacement ==> Jamie

- IMC mode matching ==> Jamie/Larisa 

- Mechanical shutters everywhere

- SRM OPLEV Connection

- MC OAF

- Better LSC whitening boards

- DAFI 

Quote:
 Anything else ?

  4358   Fri Feb 25 14:35:06 2011 Larisa ThorneUpdateElectronicsTotal harmonic distortion results for +7dBm mixer

 This experiment deals with measuring the total harmonic distortion (THD) contribution of mixers in a circuit.

(a circuit diagram is attached) where:

Mixer: ZFM-3-S+ at +7dBm

Attenuator: VAT-7+, at +7dB

Low-pass filter: SLP-1.9+, which is set to DC-1.9MHz

 

The total harmonic distortion can be calculated by the equation: 


\mbox{THD} =  \frac{V_2^2 + V_3^2 + V_4^2 + \cdots + V_\infty^2}{V_1^2}

where Vn represents the voltage of the signal at a certain harmonic n.

 

In this experiment, only the voltages of the first three harmonics were measured, with the first harmonic at 400Hz, the second at 800Hz, and the third at 1.2kHz. The corresponding voltages were read off the spectrum analyzer after it had time averaged 16 measurements. (picture of the general shape of the spectrum analyzer output is attached)


(results for this mixer's particular configuration are on the pdf attached)

 

There really isn't that much correlation between the modulations and the resulting THD.

We won't know how good these numbers are until more experiments on other mixers are done, so they can be compared. Since the rest of the mixers are relatively high levels (+17dBm, +23dBm in comparison to this experiment's +7dBm), an RF amplifier will need to be hooked up first to do any further measurements.

 

 

 

 

 

Attachment 1: THDcircuit.jpg
THDcircuit.jpg
Attachment 2: Photo_on_2011-01-17_at_12.25.png
Photo_on_2011-01-17_at_12.25.png
Attachment 3: THDwithoutamp.pdf
THDwithoutamp.pdf
  4357   Fri Feb 25 13:28:14 2011 kiwamuSummaryGeneralto do list
 Because it maybe useful to share this kind of information among us, I just put a to-do-list which has been continuously noted in my laptop.
 I guess those items can be applied for the coming two months (i.e. March and April). Anything else ?

-- Alignment

     - C1ASS

           * finalize the model and make user friendly scripts and medm screens
           * adjustment of phase rotations
           * cavity auto alignment for green

     - MC WFS 

          * put sensing matrices
          * whitening ?

    - OPLEVs

        * binary output for OPLEV whitening
        * calibration of OPLEVs  [rad/count]
        * characterization of OPLEV servos
        * bigger trans impedance gains to have reasonable readouts


-- Optimization of Suspended Optics

  - diagonalizations

         * input matrices
         * f2p with higher precision

  - damping control

        * fix ETMY damping
         * Bounce Roll notches
         * Q adjustment for damping

  - upgrading of electronics

        * bigger trans impedance gain for OSEMs
        * widen the voltage range of AA filter boards
   

   - weekly check

       * a routine script for measuring free swinging spectra


-- Input Optics

   - PMC

         * mode matching
         * epics LO HI values

   - FSS and ISS

        * recover FSS
        * make ISS working

   - EOM

       * AM minimization
       * triple resonant box

   - doubling and RFPD for green

        * mode matching to doubling crystal
        * connect RFPD_DC_MON to ADC
        * string +/-150 V and +/-15power cables from 1X1 rack to RFPDs
        * visibility check and loss investigations for the beat RFPD
        * rearrange RF amplifiers (ZLN series) for the RFPD
        * realgin Jenne's DCPD

 

-- Length Sensing and Control

   - digital system and electronics

       * characterization of RFPDs ==> SUresh/UG
       * installation of RF generation box and distribution box ==> Suresh/U
       * new LSC model and start making useful scripts (csh ? perl ? python ?)
       * binary outputs for PD whitening
       * make item lists for ordering (?)
       * draw cool diagrams for RF cable distribution and map of LSC rack.


-- Green Locking

    - X end station

           * eliminate undesired multiple spots on RFPD
           *  connect REFLPD_DCMON to ADC
           *  remote local boost
           * demodulation phase adjustment
           *  look for a high voltage amp. (bipoler)
           * installation of a mechanical shutter
           *  ETMX_TRANS CCD camera
           * analog low pass filter for temperature control

    - Y arm green locking ==> Suresh/Bryan

        * item lists
        * preparation of base mounts
        * mode matching estimation
        * mode measurement of input beam from Lightwave
        * temeperature scanning for beat location finding with IR beam
        * Installation
       * modification of PDH box

    - digital control systems

           * user friendly medm screens
           * apply proper filters for AC-DC whitened signal blending
           * add MC2 feedback path

      - noise budget

          * in-loop and out-of-loop evaluation
          * shot noise  
          * RFPD noise modeling
          * how intensity noise couples through MFD
          * electrical noise
          * frequency noise contribution from end laser and PSL
          * calibration of arm PDH signal

   - cavity scan and handing off

         * optimization of open loop transfer function for ALS
         * auto scripts   

-- misc.

   - CCD camera
        * color filters to separate IR and green (?)
   - lab laptops
        * a laptop for each end station (?)
    - dichroic TT (?)
        * large wegded and AR coated for 532nm (?)
    - epics for RS232C      
        * RS232C for doubling oven temperature

 

  4356   Fri Feb 25 10:11:56 2011 steveUpdateSAFETYhow not to

The beam of IR for doubling  is clipping on bnc cable to green beam transmitted pd.

Attachment 1: Presentation1.jpg
Presentation1.jpg
  4355   Fri Feb 25 01:48:54 2011 valeraUpdateASCmc auto alignment status

 I made several scripts to handle the mcass configuration and sensing measurements:

- The scripts and data are in the scripts/ASS directory

- The mcassUp script restores the settings for the digital lockins: oscillator gains, phases, and filters. The MC mirrors are modulated in pitch at 10, 11, 12 Hz and in yaw at 10.5, 11.5, and 12.5 Hz. The attached plot shows the comb of modulation frequencies in the MCL spectrum.

- The mcassOn and mcassOff scripts turn on and off the dither lines by ramping up and down the SUS-MC1_ASCPIT etc gains

- The senseMCdecenter script measures the response of the MCL demodulated signals to the decentering of the beam on the optics by imbalancing the coil gains by 10% which corresponds to the shift of the optic rotation point relative to the beam by 2.65 mm (75mm diameter optic) and allows calibration of the demodulated signals in mm of decentering. The order of the steps was MC1,2,3 pitch and MC1,2,3 yaw. The output of the script can be redirected to the file and analyzed in matlab. The attached plot shows the results. The plot was made using the sensemcass.m script in the same directory.

- The senseMCmirror script measures the response of the MCL demodulated signals to the mirror offsets (SUS-MC1_ASCPIT etc filter banks). The result is shown below (the sensemcass.m script makes this plot as well). There is some coupling between pitch and yaw drives so the MC coils can use some balancing - currently all gains are unity.

- The senseMCdofs scripts measures the response to the DOF excitation but I have not got to it yet.

- The next step is to invert the sensing matrix and try to center the beams on the mirrors by feeding back to optics. Note that the MC1/MC3 pitch differential and yaw common dofs are expected to have much smaller response than the other two dofs due to geometry of this tree mirror cavity. We should try to build this into the inversion.

Attachment 1: mcditherlines.pdf
mcditherlines.pdf
Attachment 2: mcdecenter.pdf
mcdecenter.pdf
Attachment 3: mcmirror.pdf
mcmirror.pdf
  4354   Thu Feb 24 21:46:30 2011 kiwamuUpdateGreen Lockinginstalled a summing box

In this past weekend I replaced a summing amplifier for the end green PDH locking by a home-made summing circuit box in order to increase the control range.

It's been working well so far.

However due to this circuit box, the demodulation phase of the PDH locking is now somewhat different from the past, so we have to readjust it at some point.

 

(background)

    At the X end station, the voltage going to the NPRO PZT had been limited up +/- 4 V because of the summing amplifier : SR560.

Therefore the laser was following the cavity motion only up to ~ +/- 4 MHz, which is not wide enough. (it's okay for night time)

So we decided to put a passive circuit instead of SR560 to have a wider range.

 

(summing box)

   We made a passive summing circuit and put it into a Pomona box.

The circuit diagram is shown below. Note that we assume the capacitance of the 1W Innolight has the same capacitance as that of the PSL Innolight (see #3640).

summing_box.png

The feedback signal from a PDH box goes into the feedback input of the circuit.

Then the signal will be low passed with the corner frequency of 200 kHz because of the combination of RC (where R is 681 Ohm and C is capacitance of the PZT).

Because of this low pass filter, we don't drive the PZT unnecessarily at high frequency.

On the other hand the modulation signal from a function generator goes into the other input and will be high passed by 50 pF mica capacitor with the corner frequency of 200 kHz.

This high pass filter will cut off noise coming from the function generator at low frequency.

In addition to it, the 50 pF capacitor gives a sufficient amount of attenuation for the modulation because we don't want have too big modulation depth.

 

Here is a plot for the expected transfer functions.

You can see that the modulation transfer function (blue curve) has non-zero phase at 216 kHz, which is our modulation frequency.

transfer_func.png
 

 

  4353   Thu Feb 24 19:59:25 2011 kiwamuUpdateGreen Lockingwhitening filter for ALS

I forgot to mention about the whitening filter for the ALS digital control system.

As usual I used a whitening filter to have a good SNR against ADC noise, but this time our whitening scheme is little bit different from the usual our systems.

I used two ADC channels for one signal and put a digital summing point  and digital filters to keep good SNR over the frequency range of interest.

It's been working fine but it's still primitive, so I will study more about how to optimize this scheme.


ACDC.png

     The diagram above shows our scheme for the signal whitening.

Basically the SNR at DC is bad when we use only a whitening filter as shown on the bottom part of the diagram because the signal is quite tiny at DC.

On the other hand if we take raw signal into ADC as 'DC path'  shown above, the SNR is better at DC but not good at intermediate frequencies (30 mHz - 1kHz).

So the idea to keep the good SNR over the frequency range of interest is to combine these 'DC path' and 'AC path' in a clever way.

     In our case, the 'DC path' signal is not as good as the 'AC path' signal above 30 mHz, so we cut off those high frequency signals by using a digital low pass filter.

In addition to it, I put a gain of 1000 in order to match the relative gain difference between 'DC path' and 'AC path'.

Then the resultant signal after the summing point keeps the good SNR with a flat transfer function up to 1 kHz. 

Quote:

Two different measurement have been performed for a test of the green locking last night.

Everything is getting better. yes. yes.

  4352   Thu Feb 24 18:21:24 2011 kiwamuUpdateGreen Lockingin-loop and out-of-loop measurements

Two different measurement have been performed for a test of the green locking last night.

Everything is getting better. yes. yes.

green_lock_setup.png

 


[ measurement 1 : IR locking]

The X arm was locked by using the IR PDH signal as usual (#4239#4268) .

The in-loop signal at from the IR path and the out-of-loop signal at from the green beat note path were measured.

measurement1.png

Let us look at the purple curve. This is an out-of-loop measurement by looking at the green beat note fluctuation.

The rms down to 0.1 Hz used to be something like 60 kHz (see here), but now it went down to approximately 2 kHz. Good.

This rms corresponds to displacement of about 260 pm of the X arm. This is barely within the line width. The line width is about 1 nm.

 

 


[ measurement 2 : green locking]

The motion of the X arm was suppressed by using the green beat signal and feeding it back to ETMX.

After engaging the ALS servo, I brought the cavity length to the resonance by changing the feedback offset from epics.

Then took the spectra of the in-loop signal at the beat path and the out-of-loop signal at the IR PDH path.

 

 Here is a time series of TRX after I brought it to the resonance.

TRX_REFL11.png

TRX was hovering around at the maximum power, which is 144 counts.

measurement2.png

Since I put one more 10:1 filter to suppress the noise around 3 Hz, the rms of the in-loop beat spectrum went to about 1 kHz, which used to be 2 kHz (see #4341).

But the out-of-loop (IR PDH signal) showed bigger noise by a factor of 2 approximately over frequency range of from 2 Hz to 2 Hz. The resultant rms is 2.7 kHz.

The rms is primarily dominated by a peak at 22 Hz (roll mode ?).

I calibrated the IR PDH signal by taking the peak to peak signal assuming the finesse of the cavity is 450 for IR. May need a cooler calibration.

  4351   Thu Feb 24 17:42:00 2011 AidanUpdateGreen Locking15% of end laser sideband power transmitted through cavity

I did a quick calculation to determine the amount of sideband transmission through the FP cavity.

The modulation frequency of the end PDH is 216kHz. The FSR of the cavity is about 3.9MHz. So the sidebands pick up about 0.17 radians extra phase on one round trip in the cavity compared to the carrier.

The ITM reflectance is r_ITM^2 = 98.5% of power, the ETM reflection is r_ETM^2 = 95%.

So the percentage of sideband power reflected from the cavity is R_SB = r_ITM*r_ETM*(exp(i*0.17) - 1)^2 / (1 - r_ETM*r_ITM exp(i*0.17) )^2 = 0.85 = 85%

So about 15% of the sideband power is transmitted through the cavity. The ratio of the sideband and carrier amplitudes at the ETM is 0.05

So, on the vertex PD, the power of the 80MHz +/-200kHz sidebands should be around sqrt(0.15)*0.05 = 0.02 = 2% of the 80MHz beatnote.

Once we get the green and IR locked to the arm again, we're going to look for the sidebands around the beatnote.

 

 

  4350   Thu Feb 24 16:47:26 2011 steveConfigurationPSL shutter is back on the PSL output

Uniblitz mechanical shutter was placed into the beam path of the PSL output with razor beam trap. The output power was 1.39W at 2.08A

It is working from the MEDM screen "old map" C1IOO_Mech_Shutter.adl

Attachment 1: P1070441.JPG
P1070441.JPG
  4349   Thu Feb 24 11:18:08 2011 steveUpdateSUSSRM sus-cables

Cheater cables   for SRM sus tied up. They were dangling aimlessly on the floor.

Attachment 1: P1070428.JPG
P1070428.JPG
  4348   Thu Feb 24 10:56:04 2011 JenneUpdateWienerFilteringOne month of H1 S5 data is now on Rossa

Just in case anyone else wants to access it, we now have 30 days of H1 S5 DARM data sitting on Rossa's harddrive.  It's in 10min segments.  This is handy because if you want to try anything, particularly Wiener Filtering, now we don't have to wait around for the data to be fetched from elsewhere.

  4347   Thu Feb 24 00:54:33 2011 KevinUpdateElectronicsCalculated Dark Noise for POX and POY

[Kevin, Rana, Koji]

I calculated the dark noise of POX and POY due to Johnson noise and voltage and current noise from the MAX4107 op-amp using nominal values for the circuit components found in their data sheets. I found that the dark noise should be approximately 15.5 nV/rtHz. The measured dark noise values are 18.35 nV/rtHz and 98.5 nV/rtHz for POX and POY respectively. The shot noise plots on the wiki have been updated to show these calculated dark noise sources.

The measured dark noise for POY is too high. I will look into the cause of this large noise. It is possible that the shot noise measurement for POY was bad so I will start by redoing the measurement.

  4346   Wed Feb 23 16:56:17 2011 Larisa ThorneUpdateVIDEOCable laying...continued

Having finished labeling the existing cables to match their new names, we (Steve, Kiwamu and Larisa) moved on to start laying new cables and labeling them according to the list.

 

Newly laid cables include: ETMXT (235'), ETMX (235'), POP (110') and MC2 (105').  All were checked by connecting a camera to a monitor and checking the clarity of the resulting image. Note that these cables were only laid, so they are not plugged in.

 

The MC2 cable needs to be ~10' longer; it won't reach to where it's supposed to. It is currently still in its place. 

The three other cables were all a lot longer than necessary.

  4345   Wed Feb 23 16:34:42 2011 valeraConfiguration pmc lens staged

I put the PMC last mode matching lens (one between the steering mirrors) on a translation stage to facilitate the PMC mode matching.

Currently 4% of incident power is reflected by the PMC. But the reflected beam does not look "very professional" on the camera to Rana - meaning there is too much TEM20 (bulls eye) mode in the reflected beam.

I locked the  PMC  on bulls eye mode and measured  the ratio of the TEM20/TEM00 in transmission to be 1.3%. Thus the PMC mode matching is ~99% and the incident beam HOM content is ~3%.

While working on the PMC I found that the source of PMC "blinking" is not the frequency control signal from MC to the laser (the MC servo was turned off) but possibly some oscillation which could be affected even by a small change of the pump current 2.10 A to 2.08 A. I showed this behaviour to Kiwamu and we decided to leave the the current at 2.08 A for now where things look stable and investigate later.

Attachment 1: PMCrefl.JPG
PMCrefl.JPG
Attachment 2: P1070438.JPG
P1070438.JPG
Attachment 3: P1070439.JPG
P1070439.JPG
  4344   Wed Feb 23 11:53:30 2011 josephbUpdateCDSUpdated mx drivers

Alex and I updated the open mx drivers from 1.3.3 to 1.3.901 (1.4 release candidate).  We downloaded the drivers from: http://open-mx.gforge.inria.fr/

We put them in /root/open-mx-1.3.901 on the fb machine (and thus get mounted by all the front ends.).  We did configure and make and make install.

We did a quick check with /opt/mx/bin/mx_info on the fb machine at this point and realized the MAC addresses and host names were all messed up, including two listings for c1iscex with different mac addresses (neither of which was c1iscex).

We then brought all the front ends mx_streams down, brought the fb down, then cleared all the peer names with mx_hostname.  We then brought the fb up, and the front end mx_stream processes.

/opt/mx/mx_info now shows a clean and correct set of hostnames and mac addresses.  Testpoints and trends are working.

  4343   Wed Feb 23 10:37:02 2011 josephbSummaryIOOMyterious data loss: FB needs investigation

Friday: 

In addition to the other fixes, Alex rebuilt the daqd process. I failed to elog this. When he rebuilt it, he needed change the symmerticom gps offset in the daqdrc file (located in /opt/rtcds/caltech/c1/target/fb). 

On Friday night, Kiwamu contacted me and let me know the frame builder had core dumped after a seg fault.  I had him temporarily disable the c1ass process (the only thing we changed that day), and then replaced Alex's rebuilt daqd code with the original daqd code and restarted it.  However, I did not change the symmetricom offset at this point.  Finally, I restarted the NDS process.  At that point testpoints and  trends seemed to be working.

Sunday:

The daqd process was restarted sometime on Sunday night (by Valera i believe).  Apparently this restart finally had the symmetricom gps offset kick in (perhaps because it was the first restart after the NDS was restarted?).  So data was being written to a future gps time.

Monday:

Kiwamu had problems with testpoints and trends and contacted me.  I tracked down the gps offset and fixed it, but the original daqd process only started once successfully, after that is was segfault, core dump non-stop. I tried Alex's rebuilt daqd (along with putting the gps offset to the correct value for it), and it worked.  Test points, trends, excitations were checked at the point and found working.

I still do not understand the underlying causes of all these segmentation faults with both the old and new daqd codes.  Alex has suggested some new open mx drivers be installed today.

Quote:

Looks like there was a mysterious loss of data overnight; since there's nothing in the elog I assume that its some kind of terrorism. I'm going to call Rolf to see if he can come in and work all night to help diagnose the issue.


 

  4342   Wed Feb 23 08:53:58 2011 SureshUpdateElectronicsRF Distribution box: Output power levels

We wish to have roughly 2 dBm of output power on each line coming out of the RF distribution box.  So I adjusted the attenuators inside the box to get this.

I also looked at why the 2x output looked so distorted and found that the input power was around 17 dBm whereas the maximum allowed (as per the datasheet of Minicircuits MK-2) is 15dBm.  So I increased the attentuation on its input line to 5dBm (up by 2dBm)  The input power levels are around 14.6dBm now  and the distortion has come down considerably.  However the net output on the 2x lines is now down to 0.7dBm.  We will have to amplify this if we need more power.

The schematic and the power output are now like this:

RF_Distribution_box_23rdFeb.jpg


  4341   Wed Feb 23 04:56:59 2011 kiwamuUpdateGreen Lockingnoise curve update

New noise spectra of the green locking have been updated.

The plot below shows the in-loop noise spectra when the beat signal was fedback to ETMX.

The rms noise integrated from 0.1 Hz to 100 Hz went down to approximately 2 kHz.

noise_suprresion.png

The red curve was taken when the beat was controlled only by a combination of some poles sand zeros on the digital filter banks. The UGF was at 40Hz.

This curve is basically the same as that Koji took few weeks ago (see here). Apparently the rms was dominated by the peaks at 16 Hz and 3 Hz.

The blue curve was taken when the same filter plus two resonant gain filters (at 16.5 Hz and 3.15 Hz) were applied. The UGF was also at 40Hz.

Due to the resonant gain filter at 16.5 Hz, the phase margin became less, and it started oscillating at the UGF as shown in the plot.

  4340   Tue Feb 22 23:40:31 2011 KojiUpdateIOOMC mode mach improvement

As per Kiwamu's request I made a light touch to the input steering and the mode matching lens.

Here V_ref and V_trans are C1:IOO-MC_RFPD_DCMON and C1:IOO-MC_TRANS_SUM, respectively.

Result: Visibility = 1 - V_ref(resonant) / V_ref(anti_reso) = 1 - 0.74 / 5.05 = 85%

What has been done:

  • Alignment of the steering mirrors before and after the last mode matching lens
       V_ref: 2.7 ==> 2.2, V_trans: 34000 ==> 39000
  • Moving of the last mode matching lens away from the MC (+ alignment of the steering mirrors)
       V_ref: 2.2 ==> 0.74, V_trans: 39000 ==> 55000
Attachment 1: IOO_MMT_110222.png
IOO_MMT_110222.png
  4339   Tue Feb 22 23:11:42 2011 valeraUpdate new medm screens: C1ASS.adl and C1MCASS.adl
Attachment 1: C1MCASS.jpeg
C1MCASS.jpeg
Attachment 2: C1ASS.jpeg
C1ASS.jpeg
  4338   Tue Feb 22 14:37:24 2011 steveUpdateSAFETYLarisa received 40m safety training

Larisa Thorne received 40m lab specific, basic safety training. She will attend P. King's Basic Laser Safety Training Session tomorrow.

 

  4337   Tue Feb 22 11:53:38 2011 steveUpdateElectronicsRF Distribution box: assembly completed

Quote:

The mechanical assembly of RF distribution box is 99% complete.  Some of the components may be bolted to the teflon base plate if needed. 

All RF cables and DC voltage supply lines have been installed and tested.  I removed the terminal block which was acting as a distribution box for the common zero voltage line.  Instead I have used the threaded holes in the body of each voltage regulator.   This allows us to keep the supply lines twisted right up to the regulator and keeps the wiring neater.  The three regulator bodies have been wired together to provide a common zero potential point. 

I did a preliminary test to see if everything is functioning. All units are functioning well.  The output power levels may need to be adjusted by changing the attenuators. 

The 2x frequency multiplier outputs are not neat sine waves.  They seem to produce some harmonics, unlike the rest of the components.

I will post the measured power output at each point tomorrow.  The RF power meter could not be found in the 40m lab.  We suspect that it has found its way back to the PSL lab.

 

 http://www.timesmicrowave.com/wireless/index.shtml  

Frank is recommending these PhaseTrack-210 as phase stable low loss rf coax cables.

  4336   Tue Feb 22 00:41:34 2011 SureshUpdateElectronicsRF Distribution box: assembly completed

The mechanical assembly of RF distribution box is 99% complete.  Some of the components may be bolted to the teflon base plate if needed. 

All RF cables and DC voltage supply lines have been installed and tested.  I removed the terminal block which was acting as a distribution box for the common zero voltage line.  Instead I have used the threaded holes in the body of each voltage regulator.   This allows us to keep the supply lines twisted right up to the regulator and keeps the wiring neater.  The three regulator bodies have been wired together to provide a common zero potential point. 

I did a preliminary test to see if everything is functioning. All units are functioning well.  The output power levels may need to be adjusted by changing the attenuators. 

The 2x frequency multiplier outputs are not neat sine waves.  They seem to produce some harmonics, unlike the rest of the components.

I will post the measured power output at each point tomorrow.  The RF power meter could not be found in the 40m lab.  We suspect that it has found its way back to the PSL lab.

 

  4335   Tue Feb 22 00:18:47 2011 valeraConfiguration c1ioo and c1ass work and related fb crashes/restarts

I have been editing and reloading the c1ioo model last two days. I have restarted the frame builder several times. After one of the restarts on Sunday evening the fb started having problems which initially showed up as dtt reporting synchronization error. This morning Kiwamu and I tried to restart the fb again and it stopped working all together. We called Joe and he fixed the fb problem by fixing the time stamps (Joe will add details to describe the fix when he sees this elog).

The following changes were made to c1ioo model:

- The angular dither lockins were added for each optics to do the beam spot centering on MC mirrors. The MCL signal is demodulated digitally at 3 pitch and 3 yaw frequencies. (The MCL signal was reconnected to the first input of the ADC interface board).

- The outputs of the lockins go through the sensing matrix, DOF filters, and control matrix to the MC1,2,3 SUS-MC1(2,3)_ASCPIT(YAW) filter inputs where they sum with dither signals (CLOCK output of the oscillators).

- The MCL_TEST_FILT was removed

The arm cavity dither alignment (c1ass) status:

- The demodulated signals were minimized by moving the ETMX/ITMX optic biases and simultaneously keeping the arm buildup (TRX) high by using the BS and PZT2. The minimization of the TRX demodulated signals has not been successful for some reason.

- The next step is to close the servo loops REFL11I demodulated signals -> TMs and TRX demodulated signals -> combination of BS and PZTs.

The MC dither alignment (c1ioo) status:

- The demodulated signals were obtained and sensing matrix (MCs -> lockin outputs) was measured for pitch dof.

- The inversion of the matrix is in progress.

- The additional c1ass and c1ioo medm screens and up and down scripts are being made.

  4334   Mon Feb 21 23:00:06 2011 ZachSummaryelogrestarted

 again

  4333   Mon Feb 21 17:29:57 2011 ranaSummaryIOOMyterious data loss: FB needs investigation

Looks like there was a mysterious loss of data overnight; since there's nothing in the elog I assume that its some kind of terrorism. I'm going to call Rolf to see if he can come in and work all night to help diagnose the issue.

Untitled.png

  4332   Mon Feb 21 11:05:51 2011 ZachUpdateelogrestarted

 I restarted the elog using the script.

  4331   Sun Feb 20 21:22:33 2011 rana, kiwamu, valeraConfigurationIOOMC Servo Change

For some reason, Kiwamu forced us to change the MC servo electronics today. We are now combining it with the FSS box.

The MC Servo by itself was locking by just driving the NPRO PZT. Becuase of the ~30 kHz mechanical resonances of that system, our badnwidth is limited. To get higher bandwidth, we can either use a wideband frequency shifter like the AOM or just use the ole FSS combo of PZT/EOM. The old MC servo was able to get 100 kHz because it used the AOM.

So we decided to try going through the FSS box. The MC servo board's FAST output now goes into the IN1 port (500 Ohm input impedance) of the TTFSS box. This allows us to use the FSS as a kind of crossover network driving the PZT/EOM combo.

At first it didn't work because of the 5V offset that Jenne, Larisa, Koji, and Suresh put into there, so I cut the wire on the board that connected the power to the summing resistor and re-installed the MC Servo board.

We also removed the old Jenne-SURF 3.7 MHz LP between the MC mixer and servo. Also removed the Kevin-box (1.6:40) stuck onto the NPRO PZT.

We have yet to measure the UGF, but it seems OK. The PCDRIVE is too high (~5-6V) so there is still some high frequency oscillation. Needs some investigation.

* To get the FSS SLOW servo to work (change NPRO temperature to minimize PZT drive onto NPRO) I set the setpoint to 5V in the script so that we operate the FSS box output at 5V mean. I set the threshold channel to point to MC_TRANS_SUM instead of RC_TRANSPD. I also had to fix the crontab on op340m so that it would point to the right scripto_cron script which runs the FSSSlowServo, RCThermalPID.pl, etc. I also had to fix scripto_cron itself since it had the old path definitions and was not loading up the EpicsTools.pm library.

** Also, I was flabbergasted by the dog clamping on the last turning mirror into the MC. Barely touching the mount changes the alignment.

  4330   Sat Feb 19 05:25:20 2011 SureshUpdateElectronicsRF: Distribution box

Most of the RF cables required for the box are done.   There are two remaining and we will attend to these tonight. 

We expect to have finished the mechanical assembly by Sunday and start a quality test on Monday.

 

 

  4329   Sat Feb 19 01:58:20 2011 ranaSummaryElectronicsLow Noise BJT Pre-Amp

Frank put his low noise preamp info here.

I suggest that we build these (using Altium) but replace the cheapo transistors with the high class MAT03 matched BJT pair from Analog Devices.

This will allow us to have a pre-amp better matched to the noise of the mixers down to low frequency.

  4328   Fri Feb 18 20:17:07 2011 JoonhoSummaryElectronicsIsolation of Voltage regulator

Today I was working on RF distribution box.

So far I almost finished to electronically isolate voltage regulators from the box wall by inserting mica sheet, sleeve, and washers.

 

The problem I found is the resistance between wall and the voltage regulator is order of M ohms

I checked my isolation (mica sheet and sleeve and washer) but there is no problem there.

But I found that the power switch is not completely isolated from the wall.( around 800 kohm)

and that the resistance between the regulator and the wall is smaller for the regulator closer to the power switch

and greater for the regulator less closer to it.

So I think we need to put washer or sleeve to isolate the powersitch electronically from the box wall.

Suresh or I will fix this problem

[ To Suresh, I can finish the isolation when I come tomorrow. Or you can proceed to finish isolation.]

  4327   Fri Feb 18 20:06:59 2011 kiwamuSummarySUScheck f2p function on ETMX

 The plot below shows how the f2p filters work.

At -2 min I turned on the f2p filters.

 f2p_ETMX.png

  4326   Fri Feb 18 18:46:08 2011 kiwamuSummarySUSf2p done on ETMX and ITMX

The f2p measurements are done on ETMX and ITMX with the real time lockin systems.

I don't explain what is the f2p measurement in this entry, but people who are interested in it can find some details on an old elog entry here or somewhere on DCC.

So far the resultant filters looked reasonable compared with the previous SRM f2p filters.

 

- backgrounds -

  Some times ago I found that the coils on ETMX had not  been nicely balanced, and it made a POS to angle coupling when I tried green locking (see here).

In addition to that, accuracy of A2L kind of measurement including the dithering techniques depend on how well the coils are balanced.  Therefore we need to balance the coils basically at all the suspended optics.

There used to be a script for this particular purpose, called f2praio.sh. This script does measure the imbalances and then balance the coils.

However this time I used the realtime lockin system to measure the imbalances instead of using the old f2p script.

One of the reasons using the real time system is that,  some of the ezca and tds commands for the old script don't work for some reasons.

Therefore we decided to move on to the real time system like Yuta did for the A2L measurement a couple of months ago.

The f2p measurement finally gives us parameters to generate a proper set of filters for POS and also the coil gains. We apply those filters and the gains in order to eliminate the POS to angle coupling and to balance the coils.

 

- results -

The followers are the resultant filters and coil gains.

The plots below show new f2p filters according to the measurement.

f2p_ITMX.png       f2p_ETMX.png

 

ITMX (assuming pendulum POS has f0 = 1 Hz and Q = 1)

ULPOS  fz = 1.009612   Qz = 1.009612 

URPOS fz = 1.125965   Qz = 1.125965  

LLPOS  fz = 0.873725   Qz = 0.873725    

LRPOS  fz = 0.974418   Qz = 0.974418  

C1:SUS-ITMY_ULCOIL_GAIN      -1.103044

C1:SUS-ITMY_URCOIL_GAIN      0.884970

C1:SUS-ITMY_LLCOIL_GAIN      0.950650

C1:SUS-ITMY_LRCOIL_GAIN      -1.060326

 

 

ETMX (assuming pendulum POS has f0 = 1 Hz and Q = 1)

ULPOS  fz = 1.055445   Qz = 1.055445   

URPOS  fz = 1.052735   Qz = 1.052735   

LLPOS  fz = 0.944023   Qz = 0.944023   

LRPOS  fz = 0.941600   Qz = 0.941600   

C1:SUS-ETMX_ULCOIL_GAIN      -0.887550

C1:SUS-ETMX_URCOIL_GAIN      1.106585

C1:SUS-ETMX_LLCOIL_GAIN = 1.07233

C1:SUS-ETMX_LRCOIL_GAIN      -0.931013

  

The precision of the coil gains looked something like 1% because every time I ran the measurement script, the measured imbalances fluctuated at this level.

The precision of the filter gain at DC (0.01 Hz) could be worse, because the integration cycles for the measurement are fewer than that of the coil gains done at high frequency (8.5 Hz).

Of course we can make the precisions by increasing the integration cycles and the excitation amplitudes, if we want to.

  4325   Fri Feb 18 17:52:25 2011 josephb, valeraUpdateCDSc1ass updated

We updated the c1ass model to include the BS.  We removed the dither excitation of the PZTs.  PZT control goes to epics. To do this, modified the /cvs/cds/caltech/target/c1iscaux/PZT_AI.db file.  We basically have it sum both the existing epics slider and our new output from c1ass.

More importantly we updated the color scheme.

We compiled and tested the Dolphin and RFM which work.

I should note we can't figure out why testpoints are not working properly with just this model.  Alex and Joe spent well over an hour trying to debug it to no success.  Current workaround is to add what channels you want from c1ass to the DAQ recording.  Other testpoints on other models appear to be working.

Attachment 1: c1ass_updated.png
c1ass_updated.png
  4324   Fri Feb 18 15:05:49 2011 kiwamuUpdateGreen Lockingtransfer function of angle to beat note (length)

[Koji and Kiwamu]

 We took transfer functions (TF) from the angle excitations at ETMX and ITMX to the green beat note signal (i.e. angle to length TF).

It turned out that the coupling from ETMX_PIT is quite large.

I wonder how f2p of the ETMX changes this coupling. We'll see.

 

a2l_TF.png

 

The plot above shows a set of the transfer functions from the angle excitation to the green beat note.

Note that the y-axis has not been calibrated, it is just a unit of counts/counts.

You can see that the TF from ETMX_PIT to the beat (red cruve) is larger than the others by about a factor of 10 over most of the frequency range.

This means that any PIT motions on ETMX can be coupled into the green beat signal somewhat over the wide frequency range.

It looks having a resonance at 1.5 Hz, but we don't exactly know why.

At that time the coil gains on only ITMX were tuned by applying f2p filters, but ETMX wasn't because of a technical reason coming from epics.

 

- - - - measurement conditions

  * PSL laser was locked to X arm by feeding back the IR PDH signal to MC2.

  * the green laser was locked to Xarm as usual.

  * took the green beat note signal (approximately 0 dBm) into Rana's MFD with the cable length of about  6 m.

  * the output from the MFD was connected to XARM_COARSE channel without a whitening filter.

  * excitation signal was injected into either ASC_PIT or ASC_YAW. The excitation was Gaussian noise with frequency band of 10 Hz and amplitude of 300 counts.

  * only ITMY had the f2p filters, which balance the coil gains all over the frequency.

 

  4323   Fri Feb 18 13:41:22 2011 josephbUpdateCDSCDS fixes

I talked to Alex today and had two things fixed:

First the maximum length of filter names (in the foton C1SYS.txt files in /chans) has been increased to 40, from 20.  This does not increase EPICS channel name length (which is longer than 20 anyways).

This should prevent running into the case where the model doesn't complain when compiled, but we can't load filters.

Additionally, we modified the feCodeGen.pl script in /opt/rtcds/caltech/c1/core/advLigoRTS/src/epics/util/ to correctly generate names for filters in all cases.  There was a problem where the C1 was being left off the file name when in the simulink .mdl file the filter was located in a box which had "top_names"  set.

  4322   Fri Feb 18 00:35:21 2011 kiwamuSummaryGeneralto do list for Feb. 18th

DAYTIME

  -  more precise F2P measurement and modify lockin simlink model (Kiwamu)

  -  run  C1ASS to check it (Valera)

  -  take care of CDS (Joe)

  -  MC mode matching (Jenne/Koji)

   -  RF stuff  (Suresh)

  -  mode matching for doubling crystal at PSL table (low priority)

  -  OPLEV (low priority)

NIGHITTIME

 -  update the noise spectra of green locking

 -  make noise budgets

  4321   Fri Feb 18 00:13:55 2011 kiwamuUpdateCDSRe:Daqd was rebuilt, now reverted.

THANK YOU, JOE !!! 

Quote:

As one of the trouble shooting steps for the daqd (i.e. framebuilder) I rebuilt the daqd executable.

ELOG V3.1.3-